OpenStudy (anonymous):

A flat rectangular piece of aluminum has a perimeter of 70 inches. The length is 11 inches longer than the width. Find the width.

OpenStudy (anonymous):

let width w=x in length l=x+11 in perimeter=2(x+11+x)=72 2x+11=72/2=36 x=?

OpenStudy (mertsj):

Did you read what surjithayer wrote?

OpenStudy (mertsj):

2x+11=72/2=36

OpenStudy (anonymous):

So 2x+11=36?

OpenStudy (mertsj):

Nevermind. What he wrote is incorrect.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

soy perimeter is 70 correction2x+11=70/2=35

OpenStudy (mertsj):